A battery of emf E and internal resistance r is supplying current in R.the maximum dissipation of power in R is 9 wwhen currenr is 3 amp findE andr

Grade:12

1 Answers

Saurabh Koranglekar
5 years ago
Dear​Archana

Acc. to​Maximum Power Transfer Theoremstates that themaximumamount of power will be dissipated by a load resistance if it is equal to theTheveninor Norton resistance of the network supplying power,

therefore we will consider internal resistance and load resistance same

r=R

power P = i^2 * (2r)

9 = (3*3)(2r)

r= 0.5 ohm

P = V*i

V = 3v